Two angles are complementary. The sum of the measure of the first angle and one-forth the second angle is 64.5 degree. Find the measures of the angles.
What is the measure of the smaller angle?
What is the measure of the other angle?

Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

64.5+.75x=90
.75x=25.5
x=34 degrees

Answer by richwmiller(17219)

a+b=90
a+b/4=64.5
a=90-b
a=64.5-b/4
set the a's equal
90-b=64.5-b/4
35.5=3b/4
142=3b
142/3=b
34=b smaller angle
56=a larger angle
check
56+34/4=64.5
34/4=8.5
34=34